# Production and decay of hyperons in transversely polarized electron-positron collider

### v1

@@ II.0.3 $e^{+}e^{-}\to\gamma^{*}/\psi\to\Xi^{-}(\Lambda\pi^{-})\bar{\Xi}^{+}(\overline{\Lambda}\pi^{+})$
Figure 4: Upper panel: Helicity frame defined for the $\psi\to\Xi^{-}(\Lambda\pi^{-})\bar{\Xi}^{+}(\bar{\Lambda}\pi^{+})$. Lower pannel: Helicity frame defined for $\Lambda(\bar{\Lambda})\to p\pi^{-}(\bar{p}\pi^{+})$.
@@
with the substitution of $\{1,3\}\to\{2,4\}$ for $\overline{\cal D}^{\nu}_{\Xi}$. This joint angular distribution, without explicit consideration of transverse polarization of beams, has been previously calculated in the literature. Hong et al. (2023); Chen and Ping (2019). For simplicity we demonstrate those joint angular distribution of single tag events with summation over the $\overline{\Xi}$ antihyperon spin directions:
@@ III.1 Moments analysis
The above moment analysis can be used to intuitively display the polarization information of $\mathcal{P}$ in the $e^{+}e^{-}\to\Lambda\bar{\Lambda}$ process and its transfer in the $\Lambda\bar{\Lambda}$ decay. In experiments, the observed variables corresponding to these moments are constructed using kinematic variables detected in the experiment, and the values are taken as the weight factors of each event in the distribution plot. If $P_{t}=0$, these moment distributions are trivially distributed as 0; if $P_{t}\neq 0$, they should exhibit a nontrivial distribution described by the Eqs. (III.1). In order to compare with unweighted events in the experiment, we generate toy Monte-Carlo events using the joint angular distribution formula $\mathcal{W}(\xi)$ for the $e^{+}e^{-}\to\Lambda\bar{\Lambda}$ process, with parameters set to $\Delta\Phi=0.4$ rad, $\alpha_{\psi}=0.69$, $\alpha_{+}=\alpha_{-}=0.75$ Ablikim et al. (2023a), and $P_{t}=0.5$. The $\phi$ in Fig. 5 represents the azimuthal angle distribution of $\Lambda$, showing the distribution of $d\mathcal{W}(\xi)/d\phi\sim\alpha_{\psi}P_{t}^{2}\cos(2\phi)$. The $\langle\mu_{i}\rangle(i=1,2,…,5)$ represents the nontrivial moment distributions, providing an intuitive display of the existence of beam transverse polarization.
Figure 5: Various distributions in the $e^{+}e^{-}\to\Lambda\bar{\Lambda}$ process. (a). $\Lambda$ azimuthal angle distribution, (b-f): moments distributions of $N*\langle\mu_{i}\rangle(i=1,2,..,5)$, here $N$ is the number of toy Monte-Carlo events, using $e^{+}e^{-}\to\psi(3686)\to\Lambda\overline{\Lambda}$ with $\alpha=0.69$ and $\Delta\Phi=23^{\circ}$ radian Ablikim et al. (2023a), and $P_{T}=0.5$.
@@ III.2 Statistical significance test
where $V(x_{i})$ represents the variance of the parameter. We assume that the beam polarization $P_{t}$ corresponding to the $\psi(3686)$ data sample can be determined through other processes, such as $e^{+}e^{-}\to\mu^{+}\mu^{-}$ measurements. The maximum likelihood fit selects four parameters $\alpha_{\psi},\Delta\Phi,\alpha_{i}$ and $\alpha_{+}$, and the error matrix formed by them can be calculated using the following equation:
@@
Figure 6 shows the measurement sensitivity of estimated $\Delta\Phi,~{}\alpha_{\psi},~{}\alpha_{-}$, and $A_{CP}={\alpha_{-}+\alpha_{+}\over\alpha_{-}-\alpha_{+}}$ under different data statistics, with their values set to $\Delta\Phi=0.4$ radian, $\alpha_{\psi}=0.69,~{}\alpha_{-}=0.748$ and $\alpha_{+}=-0.757$. In the absence of transverse beam polarization ($P_{t}=0$), the relative error in parameter measurement is maximized for the same data statistical quantity $N$. As the value of $P_{t}$ increases from 0.3 to 0.8, it can be observed that the measurement sensitivity of these parameters increases. In other words, the application of transverse beam polarization is advantageous for enhancing the measurement sensitivity of the parameters.
Figure 6: Sensitivity of $\Delta\Phi,~{}\alpha_{\psi},~{}\alpha_{-}$ and $A_{CP}=(\alpha_{-}+\alpha_{+})/(\alpha_{+}-\alpha_{+})$ in terms of data size $N$, using $e^{+}e^{-}\to\psi(3686)\to\Lambda\overline{\Lambda}$ with $\alpha=0.69$ and $\Delta\Phi=23^{\circ}$ Ablikim et al. (2023a), and $\alpha_{-}=0.748$ and $\alpha_{+}=-0.757$ Ablikim et al. (2019d).
@@ IV Summary and Perspectives
In summary, high-energy lepton beams naturally acquire transverse polarization in a storage ring through the mechanism of self-polarization, known as the Sokolov-Ternov effect. We investigate the impact of transversely polarized beams on the observables of hyperon production and sequential decay at an electron-positron collider, utilizing helicity amplitude analysis. Through moments analysis and maximum likelihood estimation, we characterize the statistical uncertainties that transversely polarized beams may impose on constraining CP violation parameters. As demonstrated in this paper, the sensitivity to measure the CP violation can be enhanced by employing transversely polarized electron and positron beams. In our paper we take $e^{+}e^{-}\to\gamma^{*}/\psi\to\Lambda(p\pi^{-})\bar{\Lambda}(\overline{p}\pi^{+})$ and $\Xi^{-}(\Lambda\pi^{-})\bar{\Xi}^{+}(\overline{\Lambda}\pi^{+})$ as examples, and those formalism can be easily extend to other processes like $\Sigma^{+}(p\pi^{0})\bar{\Sigma}^{-}(\overline{p}\pi^{0})$ and $\Lambda_{c}^{+}(\Lambda\pi^{+})\overline{\Lambda}_{c}(\overline{\Lambda}\pi^{-})$.This effect on other observables, such as higher-order quantum electrodynamic processes, hyperon weak radiative decays (e.g., Shi et al. (2022); Xing et al. (2023)), and hadronic vacuum polarization, is a topic of future interest.

### v2

@@ II.0.3 $e^{+}e^{-}\to\gamma^{*}/\psi\to\Xi^{-}(\Lambda\pi^{-})\bar{\Xi}^{+}(\overline{\Lambda}\pi^{+})$
Figure 4: Upper panel: Helicity frame defined for the $\psi\to\Xi^{-}(\Lambda\pi^{-})\bar{\Xi}^{+}(\bar{\Lambda}\pi^{+})$. Lower panel: Helicity frame defined for $\Lambda(\bar{\Lambda})\to p\pi^{-}(\bar{p}\pi^{+})$.
@@
with the substitution of $\{1,3\}\to\{2,4\}$ for $\overline{\cal D}^{\nu}_{\Xi}$. This joint angular distribution, without explicit consideration of transverse polarization of beams, has been previously calculated in the literature. Hong et al. (2023); Chen and Ping (2019). For simplicity we demonstrate those joint angular distributions of single tag events with summation over the $\overline{\Xi}$ antihyperon spin directions:
@@ III.1 Moments analysis
The above moment analysis can be used to intuitively display the polarization information in the $e^{+}e^{-}\to\Lambda\bar{\Lambda}$ process and its transfer in the $\Lambda$ and $\bar{\Lambda}$ decay. In experiments, the observed variables corresponding to these moments are constructed using kinematic variables detected in the experiment, and the values are taken as the weight factors of each event in the distribution plot. If $P_{T}=0$, these moment distributions are trivially flat; if $P_{T}\neq 0$, they should exhibit a nontrivial distribution described by the Eqs. (III.1). In order to compare with unweighted events in the experiment, we generate toy Monte-Carlo events using the joint angular distribution formula $\mathcal{W}(\xi)$ for the $e^{+}e^{-}\to\Lambda\bar{\Lambda}$ process, with parameters set to $\Delta\Phi=0.4$ rad, $\alpha_{\psi}=0.69$,  $\alpha_{-}=0.748$ and $\alpha_{+}=-0.757$ Ablikim et al. (2023a, 2019d), and $P_{T}=0.5$. The $\phi$ in Fig. 5 represents the azimuthal angle distribution of $\Lambda$, showing the distribution of $d\mathcal{W}(\xi)/d\phi\sim\alpha_{\psi}P_{T}^{2}\cos(2\phi)$. The $\langle\mu_{i}\rangle(i=1,2,…,5)$ represents the nontrivial moment distributions, providing an intuitive display of the existence of beam transverse polarization. As a reference, the $P_{t}=0$ cases are also presented for comparison. It can be observed that they appear as flat distributions with some statistical fluctuations.
Figure 5: Various distributions in the $e^{+}e^{-}\to\Lambda\bar{\Lambda}$ process. (a). $\Lambda$ azimuthal angle distribution, (b-f): moments distributions of $N*\langle\mu_{i}\rangle(i=1,2,..,5)$, here $N$ is the number of toy Monte-Carlo events, using $e^{+}e^{-}\to\psi(3686)\to\Lambda\overline{\Lambda}$ with $\alpha=0.69$ and $\Delta\Phi=23^{\circ}$ radian Ablikim et al. (2023a), and dashed histogram is for $P_{T}=0$, histogram for $P_{T}=0.5$.
@@ III.2 Statistical significance test
where $V(x_{i})$ represents the variance of the parameter. We assume that the beam polarization $P_{T}$ corresponding to the $\psi(3686)$ data sample can be determined through other processes, such as $e^{+}e^{-}\to\mu^{+}\mu^{-}$ measurements. The maximum likelihood fit selects four parameters $\alpha_{\psi},\Delta\Phi,\alpha_{-}$ and $\alpha_{+}$, and the error matrix formed by them can be calculated using the following equation:
@@
Figure 6 shows the measurement sensitivity of estimated $\Delta\Phi,~{}\alpha_{\psi},~{}\alpha_{-}$, and $A_{CP}={\alpha_{-}+\alpha_{+}\over\alpha_{-}-\alpha_{+}}$ under different statistical events of $e^{+}e^{-}\to\psi(3686)\to\Lambda(p\pi^{-})\bar{\Lambda}(\overline{p}\pi^{+})$. In the absence of transverse beam polarization ($P_{T}=0$), the relative error in parameter measurement is maximized for the same data statistical quantity $N$. As the value of $P_{T}$ increases from 0.3 to 0.8, it can be observed that the measurement sensitivity of these parameters increases. In other words, the application of transverse beam polarization is advantageous for enhancing the measurement sensitivity of the parameters.
Figure 6: Sensitivity of $\Delta\Phi,~{}\alpha_{\psi},~{}\alpha_{-}$ and $A_{CP}=(\alpha_{-}+\alpha_{+})/(\alpha_{-}-\alpha_{+})$ in terms of data size $N$, using $e^{+}e^{-}\to\psi(3686)\to\Lambda\overline{\Lambda}$ with $\alpha_{\psi}=0.69$, $\Delta\Phi=23^{\circ}$ Ablikim et al. (2023a), and $\alpha_{-}=0.748$ and $\alpha_{+}=-0.757$ Ablikim et al. (2023a, 2019d).
If $\Lambda(\bar{\Lambda})\to~{}p\pi^{-}(p\bar{\pi}^{+})$ decays conserve CP symmetry, their decay parameters satisfy the relation $\alpha_{-}=-\alpha_{+}$. If an experiment measures $\alpha_{-}\neq-\alpha_{+}$, it implies CP violation in Λ($\Lambda(\bar{\Lambda})$ decays. The significance test for CP asymmetry can be attributed to statistical hypothesis testing as follows. The null hypothesis is that the sum of the $\Lambda$ and $\bar{\Lambda}$ decay parameters is zero, while the alternative hypothesis is that the sum is not zero. We conduct a test using toy Monte Carlo events generated with parameters $\alpha_{\psi}=0.69$, $\Delta\Phi=23^{\circ}$ Ablikim et al. (2023a), and $\alpha_{-}=0.748$ and $\alpha_{+}=-0.757$ Ablikim et al. (2023a, 2019d) for different transverse polarizations $P_{T}=0,~{}0.5$ and $0.8$. The significance is calculated as $\sqrt{-2\ln~{}L_{0}-(-2\ln~{}L_{1})}$, where $L_{0}$ and $L_{1}$ are the log-likelihood values for the null and alternative hypotheses, respectively. Using the likelihood function defined in Eq.(67), $L_{0}$ is calculated with $\alpha_{-}=-\alpha_{+}=0.7525$, while $L_{1}$ is calculated with $\alpha_{-}=0.748$ and $\alpha_{+}=-0.757$ under different $P_{T}$ assumptions. The significance is shown in Fig. 7 as a function of the number of observed events. It can be observed that the significance benefits from the non-zero transverse polarization of the $e^{+}e^{-}$ beams.
Figure 7: Significance test for CP asymmetry in $\Lambda(\bar{\Lambda})\to~{}p\pi^{-}(p\bar{\pi}^{+})$ decays as a function of the number of observed events $N$, using toy Monte Carlo events for $e^{+}e^{-}\to\psi(3686)\to\Lambda\overline{\Lambda}$ generated with parameters $\alpha_{\psi}=0.69$, $\Delta\Phi=23^{\circ}$ Ablikim et al. (2023a), and $\alpha_{-}=0.748$ and $\alpha_{+}=-0.757$ Ablikim et al. (2023a, 2019d) for different transverse polarizations $P_{t}=0,~{}0.5$ and $0.8$.
@@ IV Summary and Perspectives
In summary, high-energy lepton beams naturally acquire transverse polarization in a storage ring through the mechanism of self-polarization, known as the Sokolov-Ternov effect. We investigate the impact of transversely polarized beams on the observables of hyperon production and sequential decay at an electron-positron collider, utilizing helicity amplitude analysis. It is shown that the transverse polarization of the beams introduces a new azimuthal modulation of the events in terms of the azimuthal angle of hyperons in the laboratory frame. Through moments analysis and maximum likelihood estimation, we characterize the statistical uncertainties that transversely polarized beams may impose on constraining CP violation parameters. As demonstrated in this paper, the sensitivity to measure the hyperon decay parameter and CP violation can be enhanced by employing transversely polarized electron and positron beams. In our study, we consider $e^{+}e^{-}\to\gamma^{*}/\psi\to\Lambda(p\pi^{-})\bar{\Lambda}(\overline{p}\pi^{+})$ and $\Xi^{-}(\Lambda\pi^{-})\bar{\Xi}^{+}(\overline{\Lambda}\pi^{+})$ as examples. However, the sensitivity remains largely unchanged if the transverse polarization of beam is small, e.g. in the case of $J/\psi$ decay at BESIII. Therefore, previous measurements of $J/\psi$ decays are not affected even if the transverse polarization of beams is considered. On the other hand, at the energies of big polarization degree, e.g. $\psi(3686)$, we recommend the inclusion of this effect in future data analyses at $e^{+}e^{-}$ circular colliders, and suggest extending the formalism in this paper to other processes such as $e^{+}e^{-}\to\Sigma^{+}(p\pi^{0})\bar{\Sigma}^{-}(\overline{p}\pi^{0})$, $\Lambda_{c}^{+}(\Lambda\pi^{+})\overline{\Lambda}_{c}(\overline{\Lambda}\pi^{-})$ and $\Omega^{-}(\Lambda K^{-})\overline{\Omega}^{+}(\overline{\Lambda}K^{+})$ Zhang and Song (2023); Zhang et al. (2024) for comprehensive investigations. This effect on other observables, such as higher-order quantum electrodynamic processes, hyperon weak radiative decays (e.g., Shi et al. (2022); Xing et al. (2023)), and hadronic vacuum polarization, is a topic of future interest as well.